OpenStudy (anonymous):

I work on a ploblem by Frobenius method and I encounter a equation simple ∑_(n=0)^∞▒〖(n+s)(n+s-1)a_n x〗^(n+s+e) +∑_(n=0)^∞▒〖a_n x^(n+s) 〗=0 that e is a number that it change in different problem. I coudent solve it. please help me.

OpenStudy (anonymous):

I work on a ploblem by Frobenius method and I encounter a equation simple\[\sum_{n=0}^{\infty}(n+s)(n+s-1)a_{n}x^{n+s+e}+\sum_{n=0}^{\infty}a_{n}x^{n+s}=0\] that e is a number that it change in different problem. I coudent solve it. please help me.